# Optimization via a Control-Centric Framework

### v2

@@ Abstract
Optimization plays a central role in intelligent systems and cyber-physical technologies, where the speed and reliability of convergence directly impact performance. In control theory, optimization-centric methods are standard: controllers are designed by repeatedly solving optimization problems, as in linear quadratic regulation, $H_{\infty}$ control, and model predictive control. In contrast, this paper develops a control-centric framework for optimization itself, where algorithms are constructed directly from Lyapunov stability principles rather than being proposed first and analyzed afterward. A key element is the stationarity vector, which encodes first-order optimality conditions and enables Lyapunov-based convergence analysis. By pairing a Lyapunov function with a selectable decay law, we obtain continuous-time dynamics with guaranteed exponential, finite-time, fixed-time, or prescribed-time convergence. Within this framework, we introduce three feedback realizations of increasing restrictiveness: the Hessian-gradient, Newton, and gradient dynamics. Each realization shapes the decay of the stationarity vector to achieve the desired rate. These constructions unify unconstrained optimization, extend naturally to constrained problems via Lyapunov-consistent primal-dual dynamics, and broaden the results for minimax and generalized Nash equilibrium seeking problems beyond exponential stability. The framework provides systematic design tools for optimization algorithms in control and game-theoretic problems.
@@ I Introduction
Optimization is a cornerstone of intelligent systems and cyber-physical technologies, where it underlies decision-making, resource allocation, and learning-based control. In modern control theory, optimization-based approaches are likewise central. Prominent examples include the linear quadratic regulator (LQR) [1], $H_{2}$ and $H_{\infty}$ control [2], and model predictive control (MPC) [3], where the control design is obtained by formulating and solving an optimization problem at each step of the process. These methods may be viewed as optimization-centric frameworks for control, since the optimization problem itself constitutes the foundation upon which control laws are built.
@@
In contrast to optimization-centric approaches to control, this paper focuses on a control-centric framework for optimization itself, where continuous-time dynamics are designed as optimizers by directly invoking Lyapunov stability theory and its modern extensions. This perspective parallels the classical theory of control Lyapunov functions (CLFs), in which stabilizing feedback laws are constructed directly from the Lyapunov function. Foundational results include Artstein’s equivalence between CLF existence and feedback stabilizability [4], Sontag’s universal construction of continuous stabilizing feedbacks [5, 6], and Brockett’s necessary condition for continuous static stabilization [7]. Building on this foundation, we introduce the notion of optimization Lyapunov functions (OLFs), which transfer these control-theoretic principles to the design of optimization algorithms. In particular, tools such as finite-time (FT) stability [8], fixed-time (FxT) stability [9], and prescribed-time (PT) stability [10] provide systematic ways to guarantee convergence with user-specified temporal properties. This control-centric viewpoint enables the construction of optimizer dynamics that meet explicit convergence requirements, rather than analyzing convergence only after an algorithm has been specified.
@@
A substantial body of prior work has applied Lyapunov-based methods to study continuous-time optimization, but almost always in the form of analysis rather than design. Classical contributions include the pioneering work on saddle-point dynamics for constrained optimization in [11], and the primal-dual dynamics in [12]. The stability of saddle-point dynamics has since been analyzed under different convexity and monotonicity conditions [13]. For unconstrained optimization, normalized and signed gradient dynamics with FT convergence were introduced by Cortés [14]. Subsequent extensions investigated projection-based dynamics for constrained convex problems [15], as well as FT and FxT algorithms for more general problem classes [16, 17, 18]. For minimax problems and generalized Nash equilibrium (GNE) seeking, recent works have established FT or FxT convergence properties under appropriate assumptions [19, 20, 21]. In parallel, Ross [22, 23] showed how optimizer dynamics, including coordinate descent methods, can be systematically derived from optimal control principles, highlighting the long-standing interplay between control theory and optimization.
@@ II-A Stationarity Vector
The central object in our framework is the stationarity vector, $\mathbf{S}(\mathbf{z}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, which encodes the stationarity conditions of the optimization problem at hand. The form of $\mathbf{S}$ depends on the problem addressed (unconstrained, constrained, minimax, or GNE), where the vector $\mathbf{z}$ encodes the relevant variables, which can be both the primal and dual variables. The unifying principle is that $\mathbf{S}(\mathbf{z}^{\star})=\mathbf{0}$ at any stationary point $\mathbf{z}^{\star}$.
@@ II-B Desired Convergence Properties
Lemma II.4 (Exponential Stability [24]).
@@
Lemma II.5 (FT Stability [8]).
@@
Lemma II.6 (FxT Stability [9]).
@@
Lemma II.7 (PT Stability [10]).
@@
This work focuses on these four convergence regimes, but we note that other options exist for which the generalization of our framework to them is straightforward. For other convergence regimes, the reader is referred to [25].
@@ III-A Optimization Lyapunov Function (OLF)
A central ingredient of our framework is the OLF. Analogously to Lyapunov functions for stability, an OLF certifies convergence of a dynamical optimizer to the stationarity set. This notion allows us to treat algorithm design as a control problem: the OLF encodes the desired optimality conditions, and the choice of decay law shapes the rate of convergence.
@@ III-B Nonsingularity Condition
$\mathbf{S}(\mathbf{z})\;\notin\;\ker\big(\nabla\mathbf{S}(\mathbf{z})\big).$ (14)
@@ III-C Feedback Designs Enforcing the Convergence Law
The term Newton Dynamics (ND) reflects that in the unconstrained case $S(\mathbf{z})=\nabla J(\mathbf{x})$ with a quadratic Lyapunov function (2), the feedback law reduces to the continuous-time Newton method $\dot{\mathbf{x}}=-\nabla^{2}J(\mathbf{x})^{-1}\nabla J(\mathbf{x})$ that is known to guarantee exponential convergence [26].
@@
The term Gradient Dynamics (GD) is chosen in analogy with the unconstrained case: when $S(\mathbf{z})=\nabla J(\mathbf{x})$ with a quadratic Lyapunov function (2), the feedback reduces to the classical gradient descent dynamics $\dot{\mathbf{x}}=-\nabla J(\mathbf{x})$ that is known to guarantee exponential convergence [26].
@@
Sontag’s classical work [6] considered nonlinear systems affine in the control and showed that, whenever a control-Lyapunov function exists, one can explicitly construct a feedback law that ensures asymptotic stabilization. The key idea is that the feedback recipe is universal: it applies to any system once a Lyapunov function is available. Our framework plays an analogous role in optimization: given an optimization Lyapunov function, the feedback designs in Lemmas III.3III.5, and III.7 provide universal constructions that not only ensure convergence but also shape its timing with exponential, FT, FxT, or PT guarantees.
The OLF provides a template for unifying algorithm design across problem classes. Given an optimization problem, one first identifies the appropriate stationarity conditions and encodes them in $V$. Next, by selecting a convergence law, a feedback law can be constructed that enforces the desired rate of decay. In the following sections, we apply this methodology to constrained optimization, minimax formulations, and GNE seeking problems.
@@ III-D Numerical Example: Unconstrained Problem
We illustrate the three feedback laws on the $n$-dimensional, smooth, strongly convex model. We consider the classical log-sum-exp plus quadratic term function [27], for any $x\in\mathbb{R}^{n}$
@@ IV-B Fischer-Burmeister Function and Its Usage for Stationarity
We encode KKT conditions associated with the inequality constraints in (39b) via the Fischer-Burmeister (FB) function [28]:
@@
The FB function is nonsmooth near the origin, so we use its smoothed version [29]
@@
for any $\varepsilon>0$. The smoothed FB mapping ensures the feasibility of the inequality constraints up to $\varepsilon>0$, yielding an $\varepsilon$-KKT solution. One may select $\varepsilon$ arbitrarily small to approach exact feasibility while preserving differentiability and good numerical properties [29]. If exact KKT is required, one may instead use the stationarity vector presented in Remark IV.2 below.
@@
with $\mathbf{g}_{+}(\mathbf{x})=\max(\mathbf{g}(\mathbf{x}),\mathbf{0})$ and $\boldsymbol{\lambda}_{-}=\max(-\boldsymbol{\lambda},\mathbf{0})$ applied componentwise. In this case, $\mathbf{S}(\mathbf{z})$ involves max operators and is therefore nonsmooth. However, the quadratic OLF $V(\mathbf{z})=\tfrac{1}{2}\|\mathbf{S}(\mathbf{z})\|^{2}$ remains continuously differentiable [30].
We also include the following classical regularity assumptions for constrained optimization [31].
@@ IV-C Results
Assumption IV.3 guarantees the existence of the multipliers [31, Thm. 11.6]. By Assumption III.2, $\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})\neq\mathbf{0}$ whenever $\mathbf{S}(\mathbf{z})\neq\mathbf{0}$, so the denominator in (46) is well-defined. Using Eq. (46), we get
@@
We demonstrate that uniqueness and nonsingularity hold under strong convexity, affine equalities, and convex inequalities.
@@ IV-D Illustrative Example: Network Utility Maximization
Consider the classical network utility maximization (NUM) problem, widely used in the study of continuous-time primal-dual dynamics [12]. Let $\mathbf{x}\in\mathbb{R}_{\geq 0}^{S}$ denote the vector of source rates, $\mathbf{R}\in\{0,1\}^{L\times S}$ the routing matrix (with $R_{ij}=1$ if source $j$ uses link $i$), and $\mathbf{c}\in\mathbb{R}^{L}_{>0}$ the vector of link capacities. Each source $j$ has a strictly concave, continuously differentiable utility $U_{j}:\mathbb{R}_{>0}\to\mathbb{R}$. The NUM problem reads
@@
Classical primal-dual dynamics for (49) exhibit asymptotic or exponential convergence under strict concavity [12]. Within our framework, the same problem admits FT, FxT, and PT guarantees, tightening classical results while preserving the problem structure.
Figure 2 confirms that the four timing laws realize the expected decay patterns: exponential trajectories converge asymptotically, while the FT and FxT designs reach zero in bounded time. The PT law enforces convergence exactly at the user-specified horizon. Table II highlights the corresponding computational cost. All four laws terminate within comparable runtimes, with the PT case being the most efficient in this experiment. These results demonstrate that our law selector carries over unchanged to a canonical constrained problem, providing FT, FxT, and PT guarantees beyond the classical exponential behavior reported in [12].
@@
Remark IV.7 (Stationarity monitoring).
@@ V Constrained Minimax Problems
Constrained minimax problems are central in robust control, adversarial machine learning, and game-theoretic models of resource allocation, where two agents with conflicting objectives interact under shared feasibility constraints. Classical minimax results [32, 13] establish asymptotic or exponential convergence for convex-concave problems with constraints. More recent works demonstrate FT and FxT convergence in unconstrained settings [20, 18, 33]. However, convergence in FT, FxT, and PT for constrained convex-concave problems has not been documented. This section addresses this gap using our control-centric framework with the HGD case.
@@ V-A Problem Statement and Existence
Standard regularity assumptions, such as LICQ in IV.3 and Slater in IV.4, are assumed. By Sion’s minimax theorem [34], a saddle point $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ exists and $\min\max=\max\min$ holds for the convex-concave case. We also use the nonsingularity requirement already stated in Assumption III.2 in the manuscript.

### v5

@@ Abstract
Optimization plays a central role in intelligent systems and cyber-physical technologies, where speed and reliability of convergence directly impact performance. In control theory, optimization-centric methods are standard: controllers are designed by repeatedly solving optimization problems, as in linear quadratic regulation, $H_{\infty}$ control, and model predictive control. In contrast, this paper develops a control-centric framework for optimization itself, where algorithms are constructed directly from Lyapunov stability principles rather than being proposed first and analyzed afterward. A key element is the stationarity vector, which encodes first-order optimality conditions and enables Lyapunov-based convergence analysis. By pairing a Lyapunov function with a selectable decay law, we obtain continuous-time dynamics with guaranteed exponential, finite-time, fixed-time, or prescribed-time convergence. Within this framework, we introduce three feedback realizations of increasing restrictiveness: the Hessian-gradient, Newton, and gradient dynamics. Each realization shapes the decay of the stationarity vector to achieve the desired rate. These constructions unify unconstrained optimization, extend naturally to constrained problems via Lyapunov-consistent primal-dual dynamics, and broaden the results for minimax and generalized Nash equilibrium seeking problems beyond exponential stability. The framework provides systematic design tools for optimization algorithms in control and game-theoretic problems.
@@ I Introduction
Optimization is a cornerstone of intelligent systems and cyber-physical technologies, where it underlies decision-making, resource allocation, and learning-based control. In modern control theory, optimization-based approaches are likewise central. Prominent examples include the linear quadratic regulator (LQR) [kalman_contributions_1960], $H_{2}$ and $H_{\infty}$ control [doyle_state-space_1989], and model predictive control (MPC) [rawlings_model_2024], where the control design is obtained by formulating and solving an optimization problem at each step of the process. These methods may be viewed as optimization-centric frameworks for control, since the optimization problem itself constitutes the foundation upon which control laws are built.
@@
In contrast to optimization-centric approaches to control, this paper focuses on a control-centric framework for optimization itself, where continuous-time dynamics are designed as optimizers by directly invoking Lyapunov stability theory and its modern extensions. This perspective parallels the classical theory of control Lyapunov functions (CLFs), in which stabilizing feedback laws are constructed directly from the Lyapunov function. Foundational results include Artstein’s equivalence between CLF existence and feedback stabilizability [artstein_stabilization_1983], Sontag’s universal construction of continuous stabilizing feedbacks [sontag_lyapunov-like_1983, sontag_universal_1989], and Brockett’s necessary condition for continuous static stabilization [brockett_asymptotic_1983]. Building on this foundation, we introduce the notion of optimization Lyapunov functions (OLFs), which transfer these control-theoretic principles to the design of optimization algorithms. In particular, tools such as finite-time (FT) stability [bhat_finite-time_2000], fixed-time (FxT) stability [polyakov_nonlinear_2012], and prescribed-time (PT) stability [song_time-varying_2017] provide systematic ways to guarantee convergence with user-specified temporal properties. This control-centric viewpoint enables the construction of optimizer dynamics that meet explicit convergence requirements, rather than analyzing convergence only after an algorithm has been specified.
@@
A substantial body of prior work has applied Lyapunov-based methods to study continuous-time optimization, but almost always in the form of analysis rather than design. Classical contributions include the pioneering work on saddle-point dynamics for constrained optimization in [arrow_studies_1958], and the primal-dual dynamics in [feijer_stability_2010]. The stability of saddle-point dynamics has since been analyzed under different convexity and monotonicity conditions [cherukuri_saddle-point_2017]. For unconstrained optimization, normalized and signed gradient dynamics with FT convergence were introduced by Cortés [cortes_finite-time_2006]. Subsequent extensions investigated projection-based dynamics for constrained convex problems [chen_convex_2018], as well as FT and FxT algorithms for more general problem classes [sanchez-torres_fixed_2014, li_fixed_2017, garg_fixed-time_2021]. For minimax problems and generalized Nash equilibrium (GNE) seeking, recent works have established FT or FxT convergence properties under appropriate assumptions [garg_fixed-time_2022, romero_finite-time_2020, you_unified_2024]. In parallel, Ross [ross_optimal_2019, ross_derivation_2023] showed how optimizer dynamics, including coordinate descent methods, can be systematically derived from optimal control principles, highlighting the long-standing interplay between control theory and optimization.
@@ II-A Stationarity Vector
The central object in our framework is the stationarity vector $\mathbf{S}(\mathbf{z})\in\mathbb{R}^{n}$, which encodes the stationarity conditions of the optimization problem at hand. The form of $\mathbf{S}$ depends on the problem addressed (unconstrained, constrained, minimax, or GNE), where the vector $\mathbf{z}\in\mathbb{R}^{n}$ encodes the relevant variables which can be both the primal and dual variables and the dimension $n$ varies depending to the problem addressed. The unifying principle is that $\mathbf{S}(\mathbf{z}^{\star})=\mathbf{0}$ at any stationary point $\mathbf{z}^{\star}$.
@@ II-B Desired Convergence Properties
Lemma II.4 (Exponential Stability [khalil_nonlinear_2002]).
@@
Lemma II.5 (FT Stability [bhat_finite-time_2000]).
@@
Lemma II.6 (FxT Stability [polyakov_nonlinear_2012]).
@@
Lemma II.7 (PT Stability [song_time-varying_2017]).
@@
This work focuses on these four convergence regimes, but we note that other options exist for which the generalization of our framework to them is straightforward. For other convergence regimes, the reader is referred to [song_prescribed-time_2023].
@@ III-A Optimization Lyapunov Function (OLF)
A central ingredient of our framework is the OLF which, analogously to Lyapunov functions for stability, certifies convergence of a dynamical optimizer to the stationarity set. This notion allows us to treat algorithm design as a control problem: the OLF encodes the desired optimality conditions, and the choice of decay law shapes the rate of convergence.
@@ III-B Nonsingularity Condition
$\mathbf{S}(\mathbf{z})\;\notin\;\ker\big(\nabla\mathbf{S}(\mathbf{z})^{\top}\big).$ (14)
@@ III-C Feedback Designs Enforcing the Convergence Law
The term Newton Dynamics (ND) reflects that in the unconstrained case $S(\mathbf{z})=\nabla J(\mathbf{x})$ with a quadratic Lyapunov function (2), the feedback law reduces to the continuous-time Newton method $\dot{\mathbf{x}}=-\nabla^{2}J(\mathbf{x})^{-1}\nabla J(\mathbf{x})$ that is known to guarantee exponential convergence [brown_effective_1989].
@@
The term Gradient Dynamics (GD) is chosen in analogy with the unconstrained case: when $S(\mathbf{z})=\nabla J(\mathbf{x})$ with a quadratic Lyapunov function (2), the feedback reduces to the classical gradient descent dynamics $\dot{\mathbf{x}}=-\nabla J(\mathbf{x})$ that is known to guarantee exponential convergence [brown_effective_1989].
@@
Sontag’s classical work [sontag_universal_1989] considered nonlinear systems affine in the control and showed that, whenever a control-Lyapunov function exists, one can explicitly construct a feedback law that ensures asymptotic stabilization. The key idea is that the feedback recipe is universal: it applies to any system once a Lyapunov function is available. Our framework plays an analogous role in optimization: given an optimization Lyapunov function, the feedback designs in Lemmas III.3III.5, and III.7 provide universal constructions that not only ensure convergence but also shape its timing with exponential, FT, FxT, or PT guarantees.
@@ III-D Numerical Example: Unconstrained Problem
We illustrate the three feedback laws on the $n$-dimensional, smooth, strongly convex model. We consider the classical log-sum-exp plus quadratic term function [castera_continuous_2024], for any $x\in\mathbb{R}^{n}$
@@
We note that the OLF provides a template for unifying algorithm design across problem classes. Given an optimization problem, one first identifies the appropriate stationarity conditions and encodes them in $V$. Next, by selecting a convergence law, a feedback law can be constructed that enforces the desired rate of decay. In the following sections, we apply this methodology to constrained optimization, minimax formulations, and GNE seeking problems.
@@ IV-B Fischer-Burmeister Function and Its Usage for Stationarity
We encode KKT conditions associated with the inequality constraints in (39b) via the Fischer-Burmeister (FB) function [fischer_special_1992]:
@@
The FB function is nonsmooth near the origin, so we use its smoothed version [liao-mcpherson_regularized_2019]
@@
for any $\varepsilon>0$. The smoothed FB mapping ensures the feasibility of the inequality constraints up to $\varepsilon>0$, yielding an $\varepsilon$-KKT solution. One may select $\varepsilon$ arbitrarily small to approach exact feasibility while preserving differentiability and good numerical properties [liao-mcpherson_regularized_2019]. If exact KKT is required, one may instead use the stationarity vector presented in Remark IV.2 below.
@@
with $\mathbf{g}_{+}(\mathbf{x})=\max(\mathbf{g}(\mathbf{x}),\mathbf{0})$ and $\boldsymbol{\lambda}_{-}=\max(-\boldsymbol{\lambda},\mathbf{0})$ applied componentwise. In this case, $\mathbf{S}(\mathbf{z})$ involves max operators and is therefore nonsmooth. However, the quadratic OLF $V(\mathbf{z})=\tfrac{1}{2}\|\mathbf{S}(\mathbf{z})\|^{2}$ remains continuously differentiable [rockafellar_dual_1973].
We also include the following classical regularity assumptions for constrained optimization [beck_introduction_2014].
@@ IV-C Results
Assumption IV.3 guarantees the existence of the multipliers [beck_introduction_2014, Thm. 11.6]. By Assumption III.2, $\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})\neq\mathbf{0}$ whenever $\mathbf{S}(\mathbf{z})\neq\mathbf{0}$, so the denominator in (46) is well-defined. Using Eq. (46), we get
@@
Next we demonstrate that uniqueness and nonsingularity hold under strong convexity, affine equalities, and convex inequalities.
@@
Remark IV.7 (On Radial Unboundedness).
Unlike standard Lyapunov-based stability analyses, the result in Corollary IV.6 does not require the Lyapunov function to be radially unbounded in $\mathbf{z}$. This is because the dynamics are expressed in feedback form (46), which directly enforces boundedness and drives the trajectories toward the unique KKT point. Under the strong convexity and convex feasibility assumptions, this point coincides with the global optimizer. Hence, global convergence follows from the closed-loop construction itself, rather than from the growth of $V$ at infinity.
@@ IV-D Illustrative Example: Network Utility Maximization
Consider the classical network utility maximization (NUM) problem, widely used in the study of continuous-time primal-dual dynamics [feijer_stability_2010]. Let $\mathbf{x}\in\mathbb{R}_{\geq 0}^{S}$ denote the vector of source rates, $\mathbf{R}\in\{0,1\}^{L\times S}$ the routing matrix (with $R_{ij}=1$ if source $j$ uses link $i$), and $\mathbf{c}\in\mathbb{R}^{L}_{>0}$ the vector of link capacities. Each source $j$ has a strictly concave, continuously differentiable utility $U_{j}:\mathbb{R}_{>0}\to\mathbb{R}$ associated with it. The NUM problem reads
@@
Classical primal-dual dynamics for (49) exhibit asymptotic or exponential convergence under strict concavity [feijer_stability_2010]. Within our framework, the same problem admits FT, FxT, and PT guarantees, tightening classical results while preserving the problem structure.
Figure 2 confirms that the four timing laws realize the expected decay patterns: exponential trajectories converge asymptotically, while the FT and FxT designs reach zero in bounded time. The PT law enforces convergence exactly at the user-specified horizon. Table II highlights the corresponding computational cost. All four laws terminate within comparable runtimes, with the PT case being the most efficient in this experiment. These results demonstrate that our law selector carries over unchanged to a canonical constrained problem, providing FT, FxT, and PT guarantees beyond the classical exponential behavior reported in [feijer_stability_2010].
@@
Remark IV.8 (Stationarity monitoring).
@@ V Constrained Minimax Problems
Constrained minimax problems are central in robust control, adversarial machine learning, and game-theoretic models of resource allocation, where two agents with conflicting objectives interact under shared feasibility constraints. Classical minimax results [kose_solutions_1956, cherukuri_saddle-point_2017] establish asymptotic or exponential convergence for convex-concave problems with constraints. More recent works demonstrate FT and FxT convergence in unconstrained settings [romero_finite-time_2020, garg_fixed-time_2021, ozaslan_exponential_2024]. However, convergence in FT, FxT, and PT for constrained convex-concave problems has not been documented. This section addresses this gap using our control-centric framework with the HGD case.
@@ V-A Problem Statement and Existence
Standard regularity assumptions, such as LICQ in IV.3 and Slater in IV.4, are assumed. By Sion’s minimax theorem [sion_general_1958], a saddle point $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ exists and $\min\max=\max\min$ holds for the convex-concave case. We also use the nonsingularity requirement already stated in Assumption III.2 of the manuscript.